OpenStudy (anonymous):

Given the function g(x) = 4x − 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(−4). The value of g(2) is smaller than the value of g(−4). The value of g(2) is the same as the value of g(−4). The values of g(2) and g(−4) cannot be compared.

OpenStudy (bahrom7893):

g(2) = 4*2-5=8-5=3 g(-4)=4*(-4)-5=-16-5=-21

OpenStudy (bahrom7893):

g(2) is larger

OpenStudy (anonymous):

g(2) is found by substituting in 2 for x into g(x) = 4x-5, likewise, g(-4) is found by substitution as well.

OpenStudy (anonymous):

thanks guys (: